OpenStudy (anonymous):

I am completely stuck on this one! Solve 3x^2-7x-11=0 Find the x intercepts of f(x)=3x^2-7x-11=0

hero (hero):

I already posted the solution for this earlier

OpenStudy (anonymous):

but I do not understand the solution you gave

OpenStudy (anonymous):

HELLLLLLLLLLLOOOOOOOOOO

hero (hero):

I'm sorry to hear that. (that statement doesn't mean anyone is making fun of you).

OpenStudy (anonymous):

yes rick

OpenStudy (anonymous):

CAN ONE OF YOU HELP ME

OpenStudy (anonymous):

rick, can you give us a minute?

OpenStudy (anonymous):

SURE

OpenStudy (anonymous):

Ok, your solution is unfamiliar to me. The plus/minus sign is confusing to me

hero (hero):

\[\frac{7+\sqrt{181}}{6}, \frac{7-\sqrt{181}}{6} \]

OpenStudy (anonymous):

do those checkmark lines go above these?

hero (hero):

It goes exactly how you see it...only above the 181

hero (hero):

That is the only way to include the two solutions with radicals. The other way is to combine them into one, but you didn't like that solution. So there's no other way to post the solution unless you use decimals, which again, you didn't like that. 181 is prime so it is irreducible. I'm sorry

OpenStudy (anonymous):

im trying to figure out the correct format of the answer

hero (hero):

What I have posted is the correct format

hero (hero):

There's no other way

OpenStudy (anonymous):

√7+1816?

hero (hero):

NO

OpenStudy (anonymous):

seven plus the square root of 181 all divided by 6.... and seven minus the square root of 181 all divided by 6

hero (hero):

(7 + √181)/(6) (7 - √181)/(6)

OpenStudy (anonymous):

and these are the x-intercepts, or the solutions?

hero (hero):

I don't know how you got that from my solution I posted Cali. What you posted looks nothing like my solution

OpenStudy (anonymous):

yes they are the intercepts. put that into your calculator and the answer will go in the x place of (x,y) and the y will be 0

OpenStudy (anonymous):

you are right Hero but these kids have to have it spelled out sometimes lol

OpenStudy (anonymous):

cali have you not used the quadratic formula yet?

OpenStudy (anonymous):

twig, GO F... URSELF

OpenStudy (anonymous):

whoa guys

OpenStudy (anonymous):

no need to talk like that. Your mommy wouldnt be very happy to see that. I was being nice and helping you out

hero (hero):

Rickboss, I hope you're not still waiting. If you have a question, post it in the "ask a question" box to the left

OpenStudy (anonymous):

TWIG, YOU AN A-HOLE. DONT TALK CRAP ABOUT ME

OpenStudy (anonymous):

i aint talkin crap at all... I'm a grown 31 yr old woman! what the heck i need to talk sh!t about you for! Your much younger... figure high school and maybe you aint learned his method yet. not a good idea to cuss someone that helps

OpenStudy (anonymous):

yeah guys  find the area and perimeter of the rectangular .Write each answer i simplest form